OpenStudy (anonymous):

^11sqrt(x^5)*sqrt(x) The ^11 is the little number 11 before the square root. I have no idea how to solve this, when I tried I got ^22sqrt(x^5)

OpenStudy (anonymous):

Here's what it looks like:\[\sqrt[11]{x^5}*\sqrt{x}\]

OpenStudy (anonymous):

Anyone? Please help!

OpenStudy (anonymous):

I'm not quite sure what to "solve" here because this is not a complete equation. Anyway, you can simplify it:\[\sqrt[11]{x^5} = x^{5/11},\;\;\;\;\sqrt{x} = x^{1/2}\]\[\implies \sqrt[11]{x^5} \cdot \sqrt{x} = x^{5/11}*x^{1/2} = x^{5/11\; +\; 1/2} = x^{21/22}\] Rules used here: \[a^{\frac{m}{n}} = \sqrt[n]{a^m}=(\sqrt [n] a)^m\]\[a^{r+s} = a^r\cdot a^s\]

OpenStudy (anonymous):

Thank you, I had finished the problem before and gotten the same answer. I just never got back onto OpenStudy! Thanks for the clarification, I should have used better terminology.